There are 20 people in a singing group. The mean age of 12 of the members is 18 years.
The mean age of the remaining 8 is 23 years. What is the mean age of the 20 members of
the singing group?
1
Expert's answer
2020-06-01T18:08:56-0400

Average age of 12 members is 18 years.

Total age of these members is "18\times 12=216"


Average age of the remaining 8 members is 23 years.

Total age of these members is "8\times 23=184"


Now we can calculate average age of the 20 members of the singing group:

"\frac{\text{Total age}}{\text{Number of members}}=\frac{216+184}{20}=\frac{400}{20}=20"


Answer: 20 years.